Just before class starts in the spring, you take a
Answer #1

(a)
m = 60 kg
h = 3.7m
Total floors = 86
Total Heigth gained = 86*3.7 =318.2 m

Gravitational potential energy gained = m*g*h = 60*9.8*3218.2 J
Gravitational potential energy gained = 1892301.6 J

(b)
Let the total energy be x.
0.25 * x = 1892301.6
x = 7569206.4 J

(c)
1 calorie = 4.2 joules
Calories burned = 7569206.4 /4.2 = 1802192 calories
Calories burned = 1802.2 kcal

(d)
Calories converted to thermal energy = 0.75* 1802.2 kcal
Calories converted to thermal energy = 1351.65 kcal